How do you solve this problem 3(7-m) is greater than or equal to 4(2-m) Please respond as soon as possible thank you!
Answer by Alan3354(69443)

How do you solve this problem 3(7-m) is greater than or equal to 4(2-m)
-------------
3(7-m) >= 4(2-m)
21-3m >= 8-4m
Add 4m
21+m >= 8
Subtract 21
m >= -13